# Geometric Approach to Light Rings in Axially Symmetric Spacetimes

### v2

@@ Abstract
Circular photon orbits have become an attractive topic in recent years. They play extremely important roles in black hole shadows, gravitational lensings, quasi-normal modes, and spacetime topological properties. The development of analytical methods for these circular orbits has also drawn extensive attention. In our recent work, Phys. Rev. D 106, L021501 (2022), a geometric approach to circular photon orbits was proposed for spherically symmetric spacetimes. In the present study, we give an extension of this geometric approach from spherically symmetric spacetimes to axially symmetric rotational spacetimes. In such a geometric approach, light rings in the equatorial plane are determined through the intrinsic curvatures in the optical geometry of Lorentz spacetime, which gives rise to a Randers-Finsler geometry for axially symmetric spacetimes. Specifically, light rings can be precisely determined by the condition of vanishing geodesic curvature, and the stability of light rings is classified through the intrinsic flag curvature in Randers-Finsler optical geometry. This geometric approach presented in this work is generally applicable to any stationary and axially symmetric spacetime, without imposing any restriction on the spacetime metric forms. Furthermore, we provide a rigorous demonstration to show that our geometric approach yields completely equivalent results with those derived from the conventional approach (based on the effective potential of photons).
@@ I Introduction
The circular photon orbits (such as photon spheres and light rings) have emerged as highly attractive topics in the explorations of black holes and other ultra-compact astrophysical objects as black hole mimickers, with their significance manifested through multiple avenues in theoretical and observational studies. Firstly, they are directly linked to the observed photon rings in black hole optical images [1, 2, 3, 4, 5, 6]. Secondly, they have dominant influences on other astrophysical observations, such as gravitational lensing in astrophysical systems [7, 8, 9, 10]. Thirdly, circular photon orbits uncover nontrivial relationships among Lyapunov exponent, chaotic motions of photons, and quasi-normal modes in gravitational perturbations [11, 12]. The Lyapunov exponent of photons perturbed from circular orbits, which characterizes the chaotic behavior of photon motions [13], has a remarkable connection with the imaginary part of quasi-normal mode frequency in the eikonal limit [11, 14, 15, 16]. Fourthly, the existence of stable circular photon orbits may become a hint for spacetime instability. Some recent works suggested that stable circular photon orbits may induce nonlinear instability in the dynamical evolution of black holes coupled with matter fields, producing notable observational signatures in gravitational waves and quasi-normal modes [17, 18, 19, 20]. Finally, the circular photon orbits may also reveal the topological properties of spacetimes [21, 22, 24, 23, 25]. A number of studies have shown that the photon spheres and light rings in different types of spacetimes (such as black hole spacetimes, horizon-less spacetimes produced by ultracompact objects, and naked singularity spacetimes) exhibit entirely different features [27, 28, 29, 30, 31, 39, 40, 32, 33, 34, 35, 36, 37, 38, 22, 24, 23, 25, 26]. The circular photon orbits in such spacetimes have a strong constraint on the event horizon, causal structure, and black hole hairs [23, 41, 42].
@@
There are several approaches for studying circular photon orbits (or other particles’ circular orbits) in the gravitational fields. The conventional approach employs the effective potential of photons moving in gravitational fields [5, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. This approach has proven highly valuable for analytically or numerically solving the photon orbit for a given specific gravitational source (with explicit spacetime metric expressions). However, it may become less convenient when investigating general features of photon orbits that are universally held for arbitrary and unspecific spacetime metrics, rather than dealing with a particular gravitational source with a known spacetime metric form. To circumvent this limitation, topological approaches to circular photon orbits have emerged [22, 24, 23, 25]. For each circular orbit, topological invariants of auxiliary vector fields can be assigned, such as the topological index of vector fields, Brouwer’s degree of mapping, and topological charge of Duan’s $\phi$ mapping. This kind of approach has inspired a number of works on circular orbits and black hole topology [53, 55, 56, 60, 61, 62, 59, 54, 57, 58, 63], providing us with new insights and pathways for investigating general characteristics of circular orbits, topological structure of spacetimes, and the nature of gravitational fields.
@@
In addition to the methods discussed above, a novel geometric approach to circular photon orbits was proposed in our recent work [64, 65], which relies on the mathematical construction of optical geometry and its intrinsic curvatures. This approach was initially developed for spherically symmetric spacetimes, where circular photon orbits in such spacetimes are described by photon spheres. The intrinsic curvatures in optical geometry are crucial to determine photon spheres and their stability. Specifically, photon spheres are characterized by vanishing geodesic curvature, while their stability is classified by the sign of Gaussian curvature. It has been demonstrated that this geometric approach yields completely equivalent results with the conventional approach based on local extrema of effective potentials [64, 65]. Our approach to photon spheres has enabled multiple applications, inspiring many important research on related topics [66, 67, 68, 69, 70, 71, 72]. For instance, the number and distribution characteristics of stable and unstable photon spheres in various types of spacetimes (including black hole spacetimes, ultra-compact objects’ spacetime, regular spacetimes, and naked singularity spacetimes) have been obtained through geometric analysis of Gaussian curvature and geodesic curvature [66, 67]. Additionally, recent works suggested that similar geometric treatments can be used to study the massive particles’ circular orbit and the massive particle surface [68, 69, 70]. Furthermore, it is also reported that the Gaussian curvature in optical geometry has a nontrivial connection with the Lyapunov exponent correspond to unstable photon spheres [71, 72]. These studies provide solid evidence that the geometric quantities of auxiliary optical geometry can pioneer new directions in the exploration of gravitational fields and particle orbits.
@@
Although our geometric approach has successfully characterized photon spheres in spherically symmetric spacetimes, its extension to more general and astrophysically relevant scenarios is necessary and compelling. There are vast majority of gravitational systems in astrophysics that are not spherically symmetric. Observationally, most of the supermassive black holes at the galaxy center typically exhibit rapid rotation, giving rise to axially symmetric spacetimes. Other ultra-compact astrophysical objects (like pulsars), may possess significant rotation. Consequently, a physically significant advancement of our geometric approach lies in its generalization to rotating spacetimes. The backgrounds and motivations of the present work are illustrated in figure 1.
@@
In this study, we give an extension of our geometric approach on circular photon orbits to stationary and axially symmetric rotational spacetimes. The generalization from spherically symmetric spacetimes to axially symmetric spacetimes requires additional mathematical concepts and techniques. Particularly, the constructed optical geometry for rotational spacetimes is no longer a Riemannian geometry, as in the spherically symmetric cases, and it becomes a Randers-Finsler geometry $dt=\sqrt{\alpha_{ij}dx^{i}dx^{j}}+\beta_{i}dx^{i}$ composed of a Riemannian part $\alpha$ and a non-Riemannian part $\beta$. In this work, it will be shown that intrinsic curvatures in Randers-Finsler geometry (geodesic curvature and flag curvature) provide a complete description of light rings in the equatorial plane. The light ring positions are determined by the vanishing Finslerian geodesic curvature condition $\kappa_{g}^{(F)}=0$. The stability of light rings exhibits a nontrivial connection with the intrinsic flag curvature in Randers-Finsler optical geometry. The positive flag curvature indicates the light ring to be stable, while the negative flag curvature implies the light ring to be unstable. This geometric description of stable and unstable light rings is fully equivalent to those based on the local maximum and local minimum of effective potentials.
@@
The structure of this work is organized in the following way. Section I outlines the research backgrounds and motivations of our work. Section II provides a concise introduction on the optical geometry of Lorentz spacetimes, which plays a crucial role in our geometric approach to circular photon orbits. Section III presents a detailed formulation of our geometric approach to light rings in axially symmetric spacetimes, in which the curvature conditions are introduced to determine light rings and their stability. Section IV demonstrates the equivalence between our geometric approach and the conventional approach based on the effective potential of photons. In section V, representative examples are selected to show the validity and applicability of our approach. The conclusions and perspectives are summarized in section VI. Furthermore, the mathematical preliminaries required in this work are introduced in appendices, including the geodesic equation, the definition of geodesic curvature, and the introduction of intrinsic flag curvature in the 2-dimensional Randers-Finsler geometry.
@@ II Optical Geometry of Stationary and Axially Symmetric Lorentz Spacetime
The section briefly introduces the concept and properties of the optical geometry for axially symmetric (rotational asymmetric) spacetimes. The optical geometry serves as a powerful tool to investigate the motions of photons in gravitational fields. In the present work, our geometric approach to circular photon orbits (especially the light rings in axially symmetric spacetimes) is implemented using the intrinsic curvatures of optical geometry.
The underlying physical interpretation of the mathematical construction of optical geometry can be regarded as a generalization of the Fermat’s principle in curved spacetimes [73, 74, 75, 76, 77]. The optical geometry proves to be a powerful construction in the gravitational lensing studies, and it can significantly reduce the computational complexity of the gravitational deflection angles for massless and massive particles [78, 79, 80, 81, 85, 86, 87, 88, 89, 92, 93, 90, 91, 82, 83, 84]. There are several equivalent ways to construct the optical geometry of stationary spacetimes [78, 94, 95]. One of the most straightforward methods to construct the optical geometry is through a continuous mapping of spacetime geometry $ds^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}$ (which is a 4-dimensional Lorentz manifold) into a low-dimensional manifold with the null constraint $d\tau^{2}=-ds^{2}=0$ imposed [78, 79, 80]
@@
The properties of optical geometry strongly depend on the symmetries of gravitational fields and spacetime metrics. For a spherically symmetric spacetime, its optical geometry gives rise to a Riemannian manifold [78, 79, 80, 81]. For a rotational and axially symmetric spacetime, the corresponding optical geometry is described by a Randers-Finsler manifold [82, 83, 84, 97, 98, 95, 96].
@@
For any stationary axially symmetric and rotational gravitational systems, we now give an explanation that the optical geometry eventually results in a Randers-Finsler manifold. Considering the standard rotational black hole spacetime metric
@@
the optical geometry can be obtained in a similar way by imposing the null constraint $d\tau^{2}=-ds^{2}=0$. Eventually, the arc length parameter (or spatial distance parameter) in optical geometry becomes
$dt$ $=$ $\sqrt{-\frac{g_{rr}}{g_{tt}}\cdot dr^{2}-\frac{g_{\theta\theta}}{g_{tt}}\cdot d\theta^{2}+\frac{g_{t\phi}^{2}-g_{tt}g_{\phi\phi}}{g_{tt}^{2}}\cdot d\phi^{2}}$ (7)
$-\frac{g_{t\phi}}{g_{tt}}\cdot d\phi,$
@@
In the present work, we mainly focus on the circular photon orbits in axially symmetric rotational spacetimes, whose optical geometry gives a Randers-Finsler geometry in expressions (7) or (8). In such Randers-Finsler optical geometry, for any continuous curve $\gamma=\gamma(\lambda)$ parameterized by $\lambda$, the arc-length (or spatial distance) af this curve is calculated through the integration
@@
Furthermore, if the affine parameter is chosen to be the arc-length parameter $\lambda=s$, then the modulus of the tangent vector for any continuous curve becomes unit, $T||_{(x,T)}^{(F)}=1$. The photon orbits, which are null geodesics in 4-dimensional spacetime geometry, becomes spatial geodesics when they are transformed into the Randers-Finsler optical geometry, with the stationary time coordinate $t$ to be the arc-length parameter in optical geometry (see expressions (8)).
In the Randers-Finsler geometry, there are several kinds of intrinsic geometric quantities that measure the geometric properties of Finsler geometry. The geodesic curvature $\kappa_{g}^{(F)}$ and flag curvature $\mathcal{K}^{(F)}_{\text{flag}}$ are two crucial intrinsic curvatures in the 2-dimensional Randers-Finsler geometry, which play significant roles in our geometric approach to light rings in rotational spacetime. The geodesic curvature is an intrinsic curvature of a continuous curve, which measures how far this curve is from being a geodesic curve in this Finsler manifold. The proper definition of geodesic curvature can be given in analogy with Riemannian geometry, which is presented in Appendix B. In particular, for any geodesics in 2-dimensional Randers-Finsler geometry, their geodesic curvature naturally vanish ($\kappa_{g}^{(F)}=0$), and the arc-length defined in (9) becomes extreme under a local variation ($\delta L_{AB}=0$). The flag curvature can be viewed as the generalization of the Gaussian curvature into Finsler geometry, which quantifies whether a 2-dimensional surface (pr 2-dimensional subsurface) is intrinsically flat or not. The sign of flag curvature provides a non-trivial constraint on the existence of conjugate points in Finsler geometry, which will be extremely helpful for determining the stability of light rings. The more detailed introductions and discussions on geodesics, geodesic curvature, and flag curvature are given in Appendices A-C.

### v3

@@ Abstract
Circular photon orbits have become an attractive topic in recent years. They play extremely important roles in black hole shadows, gravitational lensings, quasi-normal modes, and spacetime topological properties. In our recent work, Phys. Rev. D 106, L021501 (2022), a geometric approach to circular photon orbits was proposed for spherically symmetric spacetimes. In the present study, we extend this geometric approach from spherically symmetric spacetimes to axially symmetric spacetimes. In this geometric approach, light rings in the equatorial plane are determined by the intrinsic curvatures in the optical geometry of Lorentz spacetime, which gives rise to a Randers-Finsler geometry in axially symmetric cases. Specifically, light rings can be precisely determined by the vanishing of geodesic curvature, and the stability of light rings is classified using the intrinsic flag curvature in Randers-Finsler optical geometry. This geometric approach presented in this work is generally applicable to any stationary and axially symmetric spacetime, without imposing any restriction on the spacetime metric forms. Furthermore, we provide a rigorous demonstration to show that our geometric approach yields results that are completely equivalent to those derived from the conventional approach (based on the effective potential of photons).
@@ I Introduction
The circular photon orbits (e.g., photon spheres and light rings) have emerged as highly attractive topics in the investigations of black holes and other ultra-compact astrophysical objects as black hole mimickers. Their significance is manifested through multiple avenues in theoretical and observational studies. Firstly, they are directly linked to the observed photon rings in black hole optical images [1, 2, 3, 4, 5, 6, 7]. Secondly, they have dominant influences on other astrophysical observations, such as gravitational lensing in astrophysical systems [8, 9, 10, 11]. Thirdly, circular photon orbits reveal nontrivial relationships among the Lyapunov exponent, the chaotic motion of photons, and quasi-normal modes in gravitational perturbations [12, 13]. The Lyapunov exponent of photons perturbed from circular orbits, which characterizes the chaotic behavior of photon motions [14], has a remarkable connection with the imaginary part of the quasi-normal mode frequency in the eikonal limit [12, 15, 16, 17]. Fourthly, the existence of stable circular photon orbits may serve as an indicator of spacetime instability. Some recent works suggested that stable circular photon orbits may induce nonlinear instabilities in the dynamical evolution of black holes coupled with matter fields, producing notable observational signatures in gravitational waves and quasi-normal modes [18, 19, 20, 21]. Finally, the circular photon orbits may also reveal the topological properties of spacetimes [22, 23, 25, 24, 26]. Many studies have shown that the photon spheres and light rings in different classes of spacetimes (e.g., black hole spacetimes, horizon-less spacetimes produced by ultracompact objects, and naked singularity spacetimes) exhibit entirely different features [28, 29, 30, 31, 32, 40, 41, 33, 34, 35, 36, 37, 38, 39, 23, 25, 24, 26, 27]. The circular photon orbits in such spacetimes impose strong constraints on the event horizon, causal structure, and black hole hairs [24, 42, 43].
@@
There are several approaches for studying circular photon orbits (or other particles’ circular orbits) in the gravitational fields. The conventional approach employs the effective potential of photons moving in gravitational fields [5, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. This approach has proven highly valuable for analytically or numerically solving the photon orbit when a specific gravitational source is given (with explicit spacetime metric expressions). However, it may be less convenient when investigating general features of photon orbits that are universally applicable to arbitrary spacetime metrics, rather than dealing with a particular gravitational source with a known explicit metric form. To circumvent this limitation, topological approaches to circular photon orbits have emerged [23, 25, 24, 26]. For each circular orbit, topological invariants of auxiliary vector fields can be assigned, such as the topological index of vector fields, Brouwer’s degree of mapping, and topological charge of Duan’s $\phi$ mapping. This kind of approach has inspired numerous studies on circular orbits and black hole topology [54, 56, 57, 61, 62, 63, 60, 55, 58, 59, 64], offering us new insights and pathways for investigating the general characteristics of circular orbits, the topological structure of spacetimes, and the nature of gravitational fields.
@@
In addition to the methods mentioned above, we propose a novel geometric approach to circular photon orbits in our recent work [65, 66], which relies on the mathematical construction of optical geometry and its intrinsic curvatures. This approach was initially developed for spherically symmetric spacetimes, where circular photon orbits become photon spheres. The intrinsic curvatures in optical geometry are crucial to determining photon spheres and their stability. Specifically, photon spheres are identified by the vanishing of geodesic curvature, and their stability is classified according to the sign of Gaussian curvature. Moreover, it has been demonstrated that this geometric approach yields results that are completely equivalent to those from the conventional approach based on local extrema of effective potentials [65, 66]. Our approach to photon spheres has enabled multiple applications and inspired several interesting works on related topics [67, 68, 69, 70, 71, 72, 73]. For instance, the number and distribution characteristics of stable and unstable photon spheres in various classes of spacetimes (including black hole spacetimes, ultra-compact objects’ spacetime, regular spacetimes, and naked singularity spacetimes) have been obtained through the geometric analysis of Gaussian curvature and geodesic curvature [67, 68]. Additionally, recent works have suggested that similar geometric treatments can be used to study the circular orbits of massive particles and the massive particle surface [69, 70, 71]. Furthermore, it is also reported that the Gaussian curvature in optical geometry has a nontrivial connection with the Lyapunov exponent corresponding to unstable photon spheres [72, 73]. These studies provide solid evidence that the geometric quantities of auxiliary optical geometry can pioneer new directions in the study of gravitational fields and particle orbits.
@@
Although our geometric approach has successfully characterized photon spheres in spherically symmetric spacetimes, its generalization to more general astrophysically relevant scenarios is necessary and compelling. The vast majority of gravitational systems in astrophysics are not spherically symmetric. Observationally, most of the supermassive black holes at the galaxy center typically exhibit rapid rotation, giving rise to axially symmetric spacetimes. Other ultra-compact astrophysical objects, such as pulsars, may exhibit substantial rotation. Consequently, a physically significant advancement of our geometric approach lies in its extension to rotating spacetimes. The backgrounds and motivations of the present work are illustrated in Figure 1.
@@
In this study, we provide an extension of our geometric approach to circular photon orbits to stationary and axially symmetric spacetimes. The generalization from spherically symmetric spacetimes to axially symmetric spacetimes requires additional mathematical concepts and techniques. Particularly, the constructed optical geometry for rotational spacetimes is no longer a Riemannian geometry, as in the spherically symmetric cases, and it becomes a Randers-Finsler geometry $dt=\sqrt{\alpha_{ij}dx^{i}dx^{j}}+\beta_{i}dx^{i}$ composed of a Riemannian part $\alpha$ and a non-Riemannian part $\beta$. In this work, it will be shown that intrinsic curvatures in Randers-Finsler geometry (geodesic curvature and flag curvature) provide a complete description of light rings in the equatorial plane. The light ring positions are determined by the vanishing Finslerian geodesic curvature condition $\kappa_{g}^{(F)}=0$. The stability of light rings exhibits a nontrivial connection with the intrinsic flag curvature in Randers-Finsler optical geometry. The positive flag curvature indicates the light ring to be stable, while the negative flag curvature implies the light ring to be unstable. Furthermore, we can prove that this geometric description of stable and unstable light rings is fully equivalent to the conventional description based on the local maxima and local minima of effective potentials.
@@
The organization of this work is as follows. Section I outlines the research backgrounds and motivations of our work. Section II provides a concise introduction to the optical geometry of Lorentz spacetimes, which plays a crucial role in our geometric approach to circular photon orbits. Section III presents a detailed formulation of our geometric approach to light rings in axially symmetric spacetimes. The curvature conditions to determine light rings and their stability are proposed in this section. Section IV demonstrates the equivalence between our geometric approach and the conventional approach (based on the effective potential of photons). In Section V, representative examples are selected to show the validity and applicability of our approach. The conclusions and perspectives are summarized in section VI. Furthermore, the mathematical preliminaries required for this work are presented in appendices, including the geodesic equation, the definition of geodesic curvature, and an introduction to the intrinsic flag curvature in 2-dimensional Randers-Finsler geometry.
@@ II Optical Geometry of Stationary and Axially Symmetric Lorentz Spacetime
The section briefly introduces the concept and properties of the optical geometry for axially symmetric (rotational asymmetric) spacetimes. The optical geometry serves as a powerful tool to study for studying the photon motion in gravitational fields. In the present work, our geometric approach to circular photon orbits (especially the light rings in axially symmetric spacetimes) is implemented using the intrinsic curvatures of optical geometry.
The underlying physical interpretation of the mathematical construction of optical geometry can be regarded as a generalization of Fermat’s principle in curved spacetimes [74, 75, 76, 77, 78]. The optical geometry emerges as a powerful framework for gravitational lensing studies, substantially reducing the computational complexity of the gravitational deflection angles for massless and massive particles [79, 80, 81, 82, 86, 87, 88, 89, 90, 93, 94, 91, 92, 83, 84, 85]. There are several equivalent ways to construct the optical geometry of stationary spacetimes [79, 95, 96]. One of the most straightforward methods to construct the optical geometry is through a continuous mapping of spacetime geometry $ds^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}$ (which is a 4-dimensional Lorentz manifold) into a low-dimensional manifold with the null constraint $d\tau^{2}=-ds^{2}=0$ imposed [79, 80, 81]
@@
The properties of optical geometry strongly depend on the symmetries of gravitational fields and spacetime metrics. For a spherically symmetric spacetime, its optical geometry gives rise to a Riemannian manifold [79, 80, 81, 82]. For a stationary and axially symmetric spacetime, the corresponding optical geometry is described by a Randers-Finsler manifold [83, 84, 85, 98, 99, 96, 97].
@@
For any stationary and axially symmetric gravitational systems, we now explain that the optical geometry eventually results in a Randers-Finsler manifold. Considering the standard rotational spacetime metric
@@
the optical geometry can be obtained in a similar way by imposing the null constraint $d\tau^{2}=-ds^{2}=0$. Eventually, the arc-length parameter (or spatial distance parameter) in optical geometry becomes
$dt=\sqrt{-\frac{g_{rr}}{g_{tt}}\cdot dr^{2}-\frac{g_{\theta\theta}}{g_{tt}}\cdot d\theta^{2}+\frac{g_{t\phi}^{2}-g_{tt}g_{\phi\phi}}{g_{tt}^{2}}\cdot d\phi^{2}}-\frac{g_{t\phi}}{g_{tt}}\cdot d\phi,$ (7)
@@
In the present work, we mainly focus on circular photon orbits in axially symmetric spacetimes, whose optical geometry gives a Randers-Finsler geometry in expressions (7) or (8). In such Randers-Finsler optical geometry, for any continuous curve $\gamma=\gamma(\lambda)$ parameterized by $\lambda$, the arc-length (or spatial distance) of this curve is calculated through the integration
@@
Furthermore, if the affine parameter is chosen to be the arc-length parameter $\lambda=s$, then the modulus of the tangent vector for any continuous curve becomes unit, $T||_{(x,T)}^{(F)}=1$. The photon orbits, which are null geodesics in 4-dimensional spacetime geometry, become spatial geodesics when mapped to the Randers-Finsler optical geometry, with the stationary time coordinate $t$ to be the arc-length parameter in optical geometry (see expressions (8)).
In the Randers-Finsler geometry, several intrinsic geometric quantities are used to measure the geometric properties of this geometry. The geodesic curvature $\kappa_{g}^{(F)}$ and flag curvature $\mathcal{K}^{(F)}_{\text{flag}}$ are two crucial intrinsic curvatures in the 2-dimensional Randers-Finsler geometry, which play significant roles in our geometric approach to light rings in rotational spacetimes. The geodesic curvature is an intrinsic curvature of a continuous curve, which measures how far this curve is from being a geodesic curve in this Finsler manifold. Mathematically, the geodesic curvature in Randers-Finsler geometry can be rigorously defined in analogy with Riemannian geometry, and its definition is presented in Appendix B. In particular, for any geodesics in 2-dimensional Randers-Finsler geometry, their geodesic curvatures naturally vanish ($\kappa_{g}^{(F)}=0$), and the arc-length defined in (9) becomes extreme under a local variation ($\delta L_{AB}=0$). The flag curvature can be viewed as the generalization of the Gaussian curvature to Finsler geometry, which quantifies whether a 2-dimensional surface (or 2-dimensional subsurface) is intrinsically flat. The sign of flag curvature provides a non-trivial constraint on the existence of conjugate points in Finsler geometry, which will be extremely helpful for determining the stability of light rings. The more detailed introductions and discussions on geodesics, geodesic curvature, and flag curvature are given in Appendices A-C.